how do u get 6=x-y in standard form
Answer by oberobic(2304)

Standard form is Ax + By = C.
.
You have 6 = x - y
.
Add y to both sides
y + 6 = x
.
Add x to both sides
x + y + 6 = 0
.
Subtract 6 from both sides
x + y = -6
.
Done